### v2

# JMLR: Joint Medical LLM and Retrieval Training for Enhancing Reasoning and Professional Question Answering Capability

Abstract
With the explosive growth of medical data and the rapid development of artificial intelligence technology, precision medicine has emerged as a key to enhancing the quality and efficiency of healthcare services. In this context, Large Language Models (LLMs) play an increasingly vital role in medical knowledge acquisition and question-answering systems. To further improve the performance of these systems in the medical domain, we introduce an innovative method that jointly trains an Information Retrieval (IR) system and an LLM during the fine-tuning phase. This approach, which we call Joint Medical LLM and Retrieval Training (JMLR), is designed to overcome the challenges faced by traditional models in handling medical question-answering tasks. By employing a synchronized training mechanism, JMLR reduces the demand for computational resources and enhances the model’s ability to leverage medical knowledge for reasoning and answering questions. Our experimental results demonstrate that JMLR-13B (81.2% on Amboos, 61.3% on MedQA) outperforms models using conventional pre-training and fine-tuning Meditron-70B (76.4% on AMBOSS, 60.3% on MedQA). For models of the same 7B scale, JMLR-7B(68.7% on Amboos, 51.7% on MedQA) significantly outperforms other public models (Meditron-7B: 50.1%, 47.9%), proving its superiority in terms of cost (our training time: 37 hours, traditional method: 144 hours), efficiency, and effectiveness in medical question-answering tasks. Through this work, we provide a new and efficient knowledge enhancement tool for healthcare, demonstrating the great potential of integrating IR and LLM training in precision medical information retrieval and question-answering systems.
1 Introduction
Informed clinical decision-making is fundamentally rooted in understanding and following the logical diagnosis chain. However, the expertise required to employ this diagnostic logic effectively is not universally accessible, posing a significant challenge in healthcare. Recent developments in Large Language Models (LLMs) offer potential solutions to democratize access to this essential aspect of medical knowledge (Achiam et al., 2023; Touvron et al., 2023; Singhal et al., 2022; Peng et al., 2023; Bommasani et al., 2021; Hoffmann et al., 2022; Raffel et al., 2020; Gao et al., 2020; Computer, 2023; Soldaini et al., 2023; Bubeck et al., 2023; Wei et al., 2022; Wang et al., 2022; Yang et al., 2023). LLMs’ historical development and evaluation predominantly focused on generalized tasks, with models pretrained from large amounts of data sourced from various online platforms. This approach, however, often resulted in a dearth of high-quality, domain-specific evidence, a shortfall particularly evident in the medical field Roziere et al. (2023). A viable strategy to expand medical knowledge is to initiate training with medical-purpose LLMs and then refine this training with targeted domain-specific datasets. This hybrid training approach, blending broad linguistic capabilities with specific medical expertise, has been effectively applied in various medical models (Yunxiang et al., 2023; Zhang et al., 2023; Toma et al., 2023; Chen et al., 2023b).
However, LLMs face the significant challenge of “hallucination”, where models generate plausible but incorrect or unverified information (Ji et al., 2023; Bang et al., 2023). Such errors raise serious concerns in healthcare, where precision and reliability are imperative (Ahmad et al., 2023). Hence, recent works proposed retrieval augmented generation (RAG): they first train a retriever to obtain relevant documents from a large corpus based on the input query and then train an LLM to generate a response based on the input query and the retrieved documents (Lewis et al., 2020; Borgeaud et al., 2022a; Cheng et al., 2023; Xiong et al., 2024). By retrieving validated, up-to-date documents, RAG identifies relevant knowledge and provides contextual grounding for LLMs, thereby alleviating hallucination issues (Shuster et al., 2021; Zhu et al., 2023). Although RAG shows high accuracy in benchmarks for open-domain question answering, its effectiveness in specific domains is yet in doubt because retrievers trained in the general domain usually perform worse than those fine-tuned in the specific domain. Fine-tuning retrievers requires document-query pairs, which may not be readily available for a specific domain. Labeling such medical pairs specifically for this purpose incurs additional time and financial costs. Moreover, RAG trains retriever and LLM separately. When training LLM with a frozen retriever, it gathers documents it deems most relevant, which may not always align with what the LLM requires to produce the correct answer.
To solve these challenges, we introduce Joint Medical LLM and Retrieval Training (JMLR), a novel approach that synchronizes the training of an LLM and retriever. We evaluate whether JMLR helps enhance reasoning and professional question-answering capability. Like RAG, JMLR enhances question answering by fetching relevant documents. This involves including each retrieved document in the initial input question before the LLM’s response attempt, thereby augmenting the input to train the LLM for answer generation. In addition to RAG, JMLR introduces a unique mechanism called LLM-Rank loss to train the retriever. This is achieved by evaluating the improvement in LLM’s performance upon including any candidate retrieved documents. We calculate the log probability of the LLM’s answer with each candidate, adopting the negative of these values as the relevance score for each candidate. When reducing this loss, the retriever is trained to prioritize candidate documents that significantly aid the LLM. As shown in Figure 1, JMLR trains LLM and retriever jointly in the same step and thus improves medical question answering, especially in scenarios requiring nuanced understanding and specific information retrieval.
To validate the effectiveness of JMLR in terms of reducing computational resource requirements and fully utilizing given knowledge, we designed three experimental setups. These experiments evaluated whether direct fine-tuning, instead of traditional pretraining, would improve JMLR’s performance. We utilized external resources, including the MIMIC-IV dataset (Johnson et al., 2020), medical textbooks, and diverse medical guidelines for knowledge expansion. Our research presents a novel approach compared to the conventional pretrain-finetune process (initial pretraining followed by finetuning on a downstream task). Our technique has demonstrated superior results, achieving an impressive 53.8% accuracy on the MedQA dataset and 70.2% on AMBOSS, outperforming the traditional method’s 49.8% on MedQA and 52.8% on AMBOSS. Furthermore, our 13B model surpasses the performance of both the Meditron 70B and ChatGPT models in the USMLE and AMBOSS datasets’ question-answering tasks. Moreover, evaluation by both GPT-4 and domain experts supports the superior performance of JMLR. We summarize our key contributions as follows:
We propose a new method 111The code, along with selected retrieval data that can be made public, is included in the supplementary material and will be made publicly accessible upon the paper’s acceptance. that integrates retrieval and LLM training together, resulting in JMLR surpassing the state-of-the-art open source model by 12%.
Our training method decreases the training time required for knowledge expansion to just 37 hours. This is a considerable reduction compared to the conventional pretrain-finetune process, where pretraining alone takes 127 hours and finetuning adds an additional 17 hours.
JMLR, compared to the public SOTA models, is capable of providing enhanced reasoning for improved explainability, and the performance is better than ChatGPT(winning rate is 0.63).
2 Related Work
Medical Large Language Model
Medical LLMs have evolved from adapting existing models like BERT using biomedical datasets (Gu et al., 2021; Lee et al., 2020), to integrating document linking (Yasunaga et al., 2022b) and knowledge graphs (Yasunaga et al., 2022a). Decoder-only architectures such as GPT and Llama were then trained with domain-specific data (Radford et al., 2018; Touvron et al., 2023; Soldaini et al., 2023; Wu et al., 2023; Gema et al., 2023; Yunxiang et al., 2023; Zhang et al., 2023; Labrak et al., 2024) or some synthetic data (Tran et al., 2023; Han et al., 2023; Kweon et al., 2023; Wang et al., 2023b). Recently, scaling up data and model parameters led to models like GatorTronGPT (Peng et al., 2023) and Clinical-Camel (Toma et al., 2023), with studies on Flan-PaLM and PaLM-2’s medical reasoning (Singhal et al., 2023a, b). Furthermore, researchers have developed MEDITRON (Chen et al., 2023b), an open-source suite of medical LLMs with 7B and 70B parameters, offering enhanced performance compared to state-of-the-art baselines and outperforming closed-source models like GPT-3.5 and Med-PaLM, while closely rivaling GPT-4 and Med-PaLM-2 in effectiveness. Nonetheless, the persistent augmentation in the model’s size has not yielded commensurate enhancements in its performance metrics. This parameter expansion fails to adequately remedy the phenomenon of model forgetting (Wang et al., 2023a; Luo et al., 2023), and it appears that the model’s parameters are insufficient for retaining extensive knowledge. In contrast to existing approaches, we combine retriever and LLM in our training process to mitigate the forgetting issue. Our work provides two versions of medical LLMs: 7B and 13B. Evaluation results indicate that our models surpass previous traditional pre-trained language models in performance, and our 13B model even outperforms the MEDITRON 70B model.
Retrieval-Augmented Language Models
In the clinical domain, the occurrence of factual inaccuracies by language models (LMs), attributed to the gap in medical knowledge, presents considerable risk, with potentially severe implications such as misdiagnoses Petroni et al. (2019); Sung et al. (2021); Yao et al. (2022); Singhal et al. (2023a). Enhancing LMs by incorporating pertinent information sourced from diverse knowledge repositories has been demonstrated as an effective method for boosting performance across multiple Natural Language Processing tasks, notably in the realm of language modeling (Min et al., 2022). The retrieval component remains static in traditional RAG (Retrieval-Augmented Generation) models, while updates are applied solely to the LLM. Specifically, the input is utilized as a query in a two-step process. Initially, a retrieval system selects a collection of documents, which are essentially sequences of tokens, from a designated corpus. Subsequently, a language model assimilates these retrieved documents, employing them as supplementary information to formulate a conclusive prediction (Yu, 2022; Izacard et al., 2022). However, merely integrating the documents retrieved into the LM does not necessarily mean that the LM can use external knowledge. Consequently, several researchers have shifted the focus towards devising methods for guiding models in applying external knowledge. For instance, Atlas (Izacard et al., 2022) enhances an encoder-decoder model in tandem with the retriever by treating documents as latent variables. RETRO (Borgeaud et al., 2022b) modifies the decoder-only architecture to include retrieved texts and undertakes the pretraining of the LM from the ground up. However, LLMs are not inherently structured to seamlessly integrate retrieved content, often leading to discrepancies or inconsistencies when blending pre-trained model responses with external information. Fusing externally retrieved information with the model’s internal knowledge base can be intricate and flawed. This sometimes results in an awkward or inconsistent amalgamation of information, where the LLM may struggle to incorporate the external data into its responses coherently. Even though some researchers have considered training retriever on unsupervised data so that it can provide documents that are well-matched with LLM, such as REALM (Guu et al., 2020), REPLUG (Shi et al., 2023) and RA-DIT (Lin et al., 2023), these retriever systems are trained independently rather than in synchronization with LLM training. Our method integrates and synchronously trains the retriever and LLM, enabling the retriever to return documents more suited to LLM, thus enhancing LLM’s performance. Additionally, we have also conducted comparisons with step-by-step training methods to prove the effectiveness of our method.
3 Materials and Methods
There is a notable shortfall in biomedical knowledge datasets designed for training instruction-compliant assistants. To address this, we have collected the dataset of biomedical knowledge resources, primarily sourced from a wide array of clinical guidelines and textbook data, for pretraining. For the fine-tuning phase, we have gathered an extensive collection of medical QA datasets, aimed at augmenting the question-answering capabilities of LLMs.
3.1 Datasets
Knowledge Sources:
Data quality is important for retrieval of JMLR and for pretraining and finetuning LM. In routine practice, clinicians typically encounter at least one patient-care-related question for every two patients they see. They often turn to guidelines like UpToDate 222www.uptodate.com. UpToDate is an expert-authored, evidence-based clinical knowledge system that helps clinicians to make point-of-care decisions. Other high-quality resources include textbooks and electronic health record (EHR) notes. Our collection of UpToDate, textbooks and MIMIC-IV notes (the only publicly available EHR notes) has been used as documents to retrieve.
The fine-tuning process is based on the pre-trained Llama model, where we leveraged datasets from two significant sources: the USMLE and AMBOSS, which are essential for preparing for the USMLE examinations. The USMLE dataset, sourced directly from the official USMLE website, included a range of sample questions for Step1, Step2CK, and Step3, released between June 2022 and March 2023. These questions represent the complex medical knowledge and ethical scenarios that medical students and practitioners are expected to navigate. Additionally, we utilized the AMBOSS question bank, a comprehensive resource widely used by medical professionals and students. This dataset provided an extensive array of Step1, Step2CK, and Step3-type questions, further enriching our training material with practical and diverse medical scenarios. The MedMCQA dataset comprises 194k multiple-choice questions from Indian medical entrance exams, covering 2.4k healthcare topics and 21 medical subjects (Pal et al., 2022). The MMLU dataset Hendrycks et al. (2020), covering 57 subjects, including STEM and social sciences, features four-option multiple-choice questions for each subject. We focused on nine subjects most pertinent to medical and clinical knowledge — high school biology, college biology, college medicine, professional medicine, medical genetics, virology, clinical knowledge, nutrition, and anatomy.
3.2 Joint Medical LLM and Retrieval Training (JMLR)
3.2.1 LLM
The original Llama model supports a maximum of 4k tokens, which poses a challenge when multiple documents are retrieved. To address this, we utilized the Shifted Sparse Attention (S2-Attn) mechanism (Chen et al., 2023a), which can be used to process long sequences by mitigating the high memory cost and slow processing time associated with standard self-attention in LLMs. S2-Attn achieves this by dividing the input into several sequence spans within the self-attention modules and then introducing a shifted pattern to facilitate information exchange between these spans. This approach efficiently extends the context length without incurring additional computational costs and demonstrates performance close to the standard attention baseline. The method, designed for efficient training, retains the standard self-attention architecture during inference, making it compatible with existing infrastructure and optimizations.
3.2.2 Retriever
For information retrieval we used the ColBERT model and the trained weight as the initial model. ColBERT utilizes BERT-based encoders to transform both queries and documents into bags of embeddings. A single BERT model is shared between the query and document encoders, but the inputs are distinguished by appending special tokens: for queries and for documents. The text is tokenized for queries into BERT-based WordPiece tokens , with the token placed after BERT’s sequence-start token . For documents, ColBERT segments a document d into its constituent tokens , to which we prepend BERT’s start token followed by our special token that indicates a document sequence. Queries shorter than a predefined number of tokens are padded with BERT’s special tokens to reach the length , otherwise truncated. The padded sequence of input tokens is then passed through BERT. The resulting embeddings for queries () and documents () are computed as follows: is normalized after being processed through a CNN layer applied to BERT’s output, and undergoes a similar process but includes a filtering step to remove certain embeddings.
3.2.3 JMLR Architecture
However, Colbert matches queries and documents based on the retriever similarity, returning scores and documents. To enable the retriever system to return documents that are more helpful for providing accurate answers to Llama, we constructed the JMLR architecture. This structure employs a rank loss, , which links the Llama and retriever systems. The updating of retriever parameters is influenced by Llama’s content quality. If the returned document reduces Llama’s loss function more than document , it indicates that document is more helpful for answering than document . In calculating the rank loss, we incorporate the scores generated by the retriever system. However, due to the large variance in scores, we normalize them to obtain . The essence of is that the retriever system’s score rankings should be updated based on the ranking. in .
For JMLR model training, the dataset comprises QA pairs from AMBOSS and USMLE, with the input being a question and the output being the correct choice, including the reasoning part. The top 7 documents with the highest scores are aggregated and fed into the model during the training process. Prior to training, we extracted the top 30 documents based on scores from the retriever. During each iteration, 30 scores are generated by the retriever, and we perform weighted random sampling based on these scores, with higher scores having a greater probability of selection. This approach ensures that each document inputted into the LLM (Large Language Model) is dynamic. The LLM will likely encounter different useful guidelines and some that are not useful. This step enhances the LLM’s robustness, helping it discern which knowledge from which documents is useful and which documents are irrelevant.
3.3 Training Details
Both our training process and the conventional finetuning approach employ the AdamW optimizer, with , , and . We implement a cosine learning rate schedule, incorporating a warmup phase that accounts for 10% of the training duration and decays the learning rate to 10% of its peak value. In alignment with the practices outlined in Llama 2-chat touvron2023b, our training employs a learning rate of , a weight decay factor of 0.1, and manages a batch size of 2. The finetuning phase spans 5 epochs for all iterations. However, we apply a distinct learning rate for ColBERT, set at . The optimization strategy for training ColBERT mirrors that used for Llama 2. Throughout these experiments, we utilize four A100 80 GB GPUs. The total training duration for our method, when applied to the MedQA and AMBOSS datasets, amounts to 37 hours. Additionally, the pretraining phase on medical guidelines takes approximately 127 hours, while the finetuning phase is approximately 17 hours on 7B Llama. This comprehensive training approach ensures the effective adaptation and optimization of the model for specific medical applications.
4 Main Results
4.1 Effective of Retrieval
Setup: For benchmarks with available public training sets, we individually apply supervised finetuning using JMLR on each set and subsequently assess their performance on the corresponding test sets. PubMedQA and MedMCQA both offer detailed reasoning traces, such as extensive answers or explanations. For MedQA, AMBOSS, and MedMCQA, we employ an alternative training set equipped with human-crafted explanations for each query. In this scenario, JMLR is trained to attach the answer directly to the explanation. Regarding MMLU-Medical, a dataset without its own training set, we opt to evaluate it using the JMLR model originally trained on MedQA and AMBOSS. We calculate the accuracy of different medical QA datasets based on different models.
| Datasets | MMLU-Medical | MedMcQA | MedQA | AMBOSS |
| PMC-Llama-7B | 59.7 | 57.6 | 42.4 | 43.7 |
| Llama 2-7B | 56.3 | 54.4 | 44.0 | 46.5 |
| Meditron-7B | 55.6 | 59.2 | 47.9 | 50.1 |
| JMLR-7B | 57.2 | 61.3 | 51.7 | 68.7 |
Results: In Table 1, our JMLR-7B model, operating at the 7B scale, demonstrates its superiority by outperforming a variety of state-of-the-art (SOTA) public models across several benchmark datasets in the medical field. The datasets include MMLU-Medical, MedMcQA, MedQA, and AMBOSS. The enhanced performance of JMLR-7B can be partially attributed to its use of Llama-2 as the foundational model, showcasing significantly higher average performance than other pre-trained baselines. However, the distinct edge comes from our integrated training approach that combines retriever and language model learning (LLM). This methodology not only further boosts the performance of Llama-2 in medical benchmarks but also enables the retriever component to effectively assist the LLM in sourcing relevant and helpful documents for answering questions. This synergy significantly reduces instances of hallucination and enhances the overall effectiveness of the model. In the MMLU-Medical and MedMcQA datasets, JMLR-7B achieved scores of 57.2% and 61.3%, respectively, surpassing the scores of PMC-Llama-7B (59.7% and 57.6%), Llama 2-7B (56.3% and 54.4%), and Meditron-7B (55.6% and 59.2%). The trend continues in the MedQA and AMBOSS datasets, where JMLR-7B not only outperforms these models but does so with a notable margin, scoring 51.7% in MedQA and an impressive 68.7% in AMBOSS. This is particularly significant when compared to Meditron-7B, which scores 47.9% and 50.1% in these datasets, respectively. Overall, the results illustrate that the JMLR-7B model, on average, achieves about a 10% improvement in performance over its closest competitor, Meditron-7B. This highlights the effectiveness of our training method and the substantial advancements it brings to the field of medical benchmarking.
4.2 Effective of Joint LLM and Retrieval Training
Setup: In our experiment, we compared the 7B and 13B Llama models. We employed conventional RAG, fine-tuning the 7B and 13B models(RAG-7B and RAG-13B). During training, we didn’t update the parameters of the retriever; only the parameters of Llama were updated. We also separated the training steps: initially, we fixed the parameters of the retriever and only updated those of Llama, and then fixed Llama’s parameters while updating the retriever’s using rank loss, resulting in JLMR separate models, JLMR-separate-7B and JLMR-separate-13B. Finally, we trained JLMR using our integrated approach to obtain JMLR-7B and JMLR-13B. Additionally, we compared these models with larger-scale models such as ChatGPT and Meditron-70B.
Results: As shown in Figure 2, while the traditional RAG method offers some model accuracy enhancement, its impact is not particularly substantial. However, when we update the retriever using our rank loss(our-separate-7B and our-separate-13B), there is a noticeable improvement over the traditional RAG. Our-separate-7B and our-separate-13B achieved 51.8% and 65.4% in MedQA, and 54.5% and 71.3% in AMBOSS, respectively, outperforming the RAG variants. This improvement stems from the retriever’s enhanced ability, after the update, to identify documents that aid Llama in accurately answering questions. However, asynchronous training can be a bottleneck, especially in the efficient use of medical guidelines. While updating Llama’s parameters, the static nature of the retriever’s guidelines means that Llama might miss out on some crucial guidelines. This gap leads to the underutilization of potentially beneficial medical guidelines. Our joint training process,(JMLR-7B and JMLR-13B), effectively mitigates this issue, leading to a significant boost in performance. Because the parameters of the retriever are continuously updated during our training method, the best documents returned also change, allowing the model to encounter different documents that are helpful in answering questions, thereby enabling it to make better use of these medical guidelines. Compared to the larger models, JMLR-13B not only outperforms ChatGPT (which scores 55.4% in MedQA and 55.6% in AMBOSS) but also surpasses Meditron-70B. Specifically, JMLR-13B achieves an impressive 61.3% in MedQA and 81.2% in AMBOSS, indicating a substantial leap in model effectiveness and underscoring the success of our joint training approach.
4.3 Effective of Diverse Training Processes
Setup: To explore the impact of pretraining, finetuning, and information retrieval on model performance, we designed the following experiments: We tested the original Llama 2-7B on MedQA and AMBOSS datasets (Llama-7B). We also added a retriever during the inference stage to extract documents to assist Llama 2-7B in answering questions(Llama-7B+Retriever). Additionally, we finetuned Llama 2-7B on different training sets to obtain FT-Llama-7B. The same operation was repeated, adding a retriever during the inference stage to extract relevant documents to assist the model in answering questions(FT-Llama-7B+Retriever). Furthermore, to investigate the effectiveness of pretraining, we pretrained Llama 2 on medical guidelines(Llama-7B-Pretrain) and repeated the tests with and without adding a retriever (FT-Llama-7B-Pretrain, FT-Llama-7B-Pretrain+Retriever). Finally, we applied our training approach for further training on Llama-7B and Llama-7B-Pretrain (JMLR-7B, JMLR-7B-Pretrain).
Result: Figure 3 displays the varied performance of models at different stages. We observed that Llama-7B+Retriever, compared to Llama 2-7B, did not improve but decreased effectiveness. This suggests that Llama 2-7B does not utilize these guidelines effectively, and longer guidelines may even lead to misunderstandings and hallucinations. In contrast, FT-Llama-7B significantly improved over Llama-7B, indicating that the finetuning phase effectively enhanced Llama’s medical QA capabilities. However, this phase did not aid in the utilization of medical guidelines. FT-Llama-7B-Pretrain further improved upon FT-Llama-7B, and adding Retriever during the inference stage resulted in a slight increase in performance. This improvement is likely due to the pretraining phase expanding Llama 2-7B’s medical knowledge, thereby enabling it to utilize medical guidelines more effectively. Lastly, JMLR-7B and JMLR-7B-Pretrain achieved the best results. However, the minor difference between JMLR-7B and JMLR-7B-Pretrain suggests that our method has already sufficiently expanded Llama 2-7B’s medical knowledge. Considering training time, JMLR-7B required only 37 hours, while JMLR-7B-Pretrain needed 127 hours (for pretraining) plus 37 hours (for our method), indicating that our approach is more efficient in expanding the model’s medical knowledge.
4.4 Effective of the reasoning ability
Setup: To verify that our model can accurately answer questions and provide detailed reasoning, we tasked JMLR-13B with generating reasons for its responses to the AMBOSS and MedQA test sets. Similarly, we had ChatGPT generate corresponding reasons. We employed two methods for scoring these reasons: 1. We used GPT-4 to judge which model’s generated reasoning was better based on the ground truth. 2. We engaged three medical professionals to review and assess a sample of 20 questions drawn from the test sets. These doctors were asked to judge which model’s reasoning was closer to the actual, clinically accurate reasons.
Result: Table 2 demonstrates the varying win rates of our model under two different evaluation methods. In the assessment using GPT-4, JMLR-13B achieved a win rate of 0.63, while under the evaluation of medical professionals, our model attained a win rate of 0.60. These results suggest that simply increasing the model’s parameters is not the most effective way to expand medical knowledge. Instead, JMLR-13B’s ability to consult different medical guidelines for various questions and to find corresponding answers and explanations significantly reduces hallucinations and greatly enhances the model’s explanatory power. This achievement highlights the model’s proficiency in accessing and applying relevant medical information contextually, thereby providing more accurate and credible explanations. The close alignment of the win rates from both GPT-4 and medical professionals’ evaluations further underscores the model’s robustness and reliability in medical question-answering and reasoning tasks.
| Comparison | Win Rate |
| GPT4 | 0.63 |
| Expertise | 0.6 |
5 Conclusion
This study introduces the Joint Medical LLM and Retrieval Training (JMLR), significantly enhancing performance in medical question-answering and reasoning tasks. The JMLR models not only outperform existing state-of-the-art models in efficiently handling medical resources but also effectively reduce hallucinations in information generation, improving the accuracy and reliability of answers and explanations. Experimental results demonstrate substantial performance improvements across multiple medical benchmark tests, validating the effectiveness of integrating retriever and language model training. Furthermore, the reasoning capabilities of the JMLR model have been recognized by both GPT-4 and medical professionals, further confirming its potential and reliability for applications in the medical domain.
6 Limitations and Ethical Considerations
This study offers valuable insights but also comes with several limitations that we would like to highlight:
Domain Specificity: Our research exclusively focuses on the task of medical QA and reasoning. The adaptation of the proposed method to other domains remains unexplored. This suggests that our approach may need further validation and adjustments before being applied to different fields.
Expertise of Annotators: We relied on 3 doctors as annotators for human evaluation and preference results. While they are qualified to provide expert medical opinions and insights, employing more qualified domain experts as annotators would enhance the statistical significance of our results. We leave this to future work, along with addressing concerns about fairness, generalizability to other domains/languages, and potential biases inherent in LLMs.
Privacy Implications
Despite significant advancements in the performance of medical knowledge acquisition and question-answering systems, privacy protection becomes a paramount concern when dealing with sensitive medical data. In particular, our model utilizes data from textbooks and Uptodate. Textbooks are public datasets, while the UpToDate dataset cannot be publicly disclosed or used to produce derivative products based on their terms of use. We have sent an email to the official authorities and received their permission to use it. We would like to express our gratitude to UpToDate and must ensure that no information about UpToDate is disclosed.
Bias Considerations
Furthermore, while we strive to enhance the model’s performance across multiple medical question-answering datasets, we must acknowledge that the choice and composition of datasets could introduce or exacerbate biases within the model. For example, if training data predominantly comes from certain geographic locations or populations, the model might exhibit biases towards medical conditions or treatment methods outside those groups. This could limit the fairness and effectiveness of the model when applied globally. Future research should consider training with more diverse and comprehensive datasets to reduce potential biases and enhance the model’s universal applicability.
Broader Impacts
Our study represents an important step forward in advancing AI applications in the field of medicine, but its broader societal impacts also require careful consideration. In particular, automated medical question-answering systems hold great potential in improving healthcare efficiency and accuracy but could also impact the roles of medical professionals and patient care practices. For instance, reliance on these systems may sometimes reduce direct communication between doctors and patients or might influence physicians’ clinical judgment. Therefore, implementing these technological solutions should be approached with caution, ensuring they serve as a complement, not a replacement, to the toolkit of medical professionals. Moreover, the public’s understanding and acceptance of these systems are crucial for their widespread use, necessitating enhanced education and transparency.
In summary, while our research demonstrates the potential of leveraging large language models and information retrieval techniques in medical question-answering systems, close attention must be paid to the ethical considerations of privacy, bias, and broader societal impacts. Future work should aim to address these challenges, ensuring the development and application of these technologies benefit the healthcare system and society as a whole.
Appendix A Appendix
A.1 Case Study
We presented four specific examples to further understand why JLMR outperforms other public models. The first three examples are from the AMBOSS dataset. Due to privacy concerns with AMBOSS, we cannot display the complete content of these questions. The last example is from a USMLE question. Our answers for the first two questions are correct, while ChatGPT responded incorrectly. Although both our model and ChatGPT answered the last two questions correctly, there were some differences in the reasoning generated. For the first example, we found that JLMR’s retrieval function could extract similar questions, leading the model to answer incorrectly regarding direction. The retrieved document served as a background, fully utilizing the previously fine-tuned dataset. In the second example, we noticed that ChatGPT made a common-sense error: vaccination is not recommended for pregnant women in such cases. Even though ChatGPT has learned far more medical documents than our model by using pretraining or finetuning, it does not mean that the model will apply all medical knowledge correctly, even for some basic medical issues. In the third example, both we and ChatGPT answered correctly, but the reasoning generated by ChatGPT differed significantly from the correct answer’s reasoning: ChatGPT simply stated some facts without providing a logical reason. However, our model gave a more detailed and logical explanation of why there are decreased circulating T cells, as JLMR could retrieve related documents, such as some documents about SCID, to better understand the underlying principles. The last example differs from the first three, as the highest-scoring document retrieved was from UpToDate’s guideline. The guideline provides a detailed introduction to PCOS and its symptoms, helping the model better explain. In contrast, ChatGPT simply correlated the symptoms straightforwardly without thoroughly analyzing other related symptoms, which can often lead to diagnostic errors, even though it answered this question correctly. Our model, however, performed a comprehensive analysis.
A.2 Accuracy Variation with Different Numbers of Documents
During training, we set the number of background documents to seven, maintaining this count during the inference phase as well. This quantity is optimal, according to our experiments, as illustrated in the figure 4. We employed the JLMR method to train the LLM on the MedQA dataset, retrieving varying numbers of documents. The performance is at its weakest when only one document is retrieved; this insufficiency leads to a lack of adequate medical knowledge for the model. As we increase the number of retrieved documents, the model’s performance gradually improves, reaching its peak with seven documents. However, once this number increases to ten, performance again declines due to retrieving an excess of irrelevant documents, which hinders the model’s ability to answer questions effectively.
| 1 | Q1: A 48-year-old man… |
| Answer: | |
ChatGPT: The correct answer is A) Cerebral edema. Explanation: This patient presents with symptoms consistent with alcohol withdrawal syndrome (AWS), which can manifest as seizures (alcohol withdrawal seizures) due to abrupt cessation or reduction of alcohol intake in a chronic heavy drinker. AWS seizures typically occur within 48 hours of reducing alcohol intake. This patient’s laboratory findings are indicative of hyponatremia (low serum sodium), which is common in severe alcohol withdrawal due to the syndrome of inappropriate antidiuretic hormone secretion (SIADH). SIADH leads to water retention and dilutional hyponatremia. The risk of cerebral edema is increased in patients with severe hyponatremia who are rapidly corrected to normal serum sodium levels. Rapid correction of hyponatremia can lead to osmotic shifts of water into the brain cells, causing cerebral edema, which may result in neurological deterioration, seizures, coma, and death. This risk is particularly high in chronic alcoholics due to adaptive changes in brain osmolytes. Given the patient’s current status, urgent treatment to normalize serum sodium levels (such as with hypertonic saline infusion) puts him at increased risk for cerebral edema. Therefore, careful monitoring and slow correction of hyponatremia are essential in this patient to avoid this potentially fatal complication. |
Our: This patient’s presentation is consistent with alcohol withdrawal syndrome (AWS), which is characterized by autonomic hyperactivity, tremulousness, and seizures. AWS can occur in patients with a history of chronic alcohol use disorder who abruptly stop drinking or significantly reduce their alcohol intake. The syndrome is thought to be due to the withdrawal of the inhibitory effects of alcohol on the central nervous system. The most common complication of AWS is seizures, which can occur within 48 hours of the last drink. Seizures are often generalized tonic-clonic seizures and can be life-threatening. Other complications include delirium tremens (DTs), which is characterized by hallucinations, confusion, and autonomic instability. DTs can also lead to seizures and death. The treatment of AWS involves supportive care, including hydration, benzodiazepines, and possibly antipsychotics. In severe cases, hospitalization may be necessary for close monitoring and management of complications. Osmotic myelinolysis (also known as central pontine myelinolysis) is a condition that can occur in patients with severe hyponatremia (low sodium levels in the blood) who are treated with rapid correction of their sodium levels. This condition is characterized by demyelination of the central nervous system, leading to neurologic symptoms such as confusion, seizures, and coma. Given the patient’s presentation and history of chronic alcohol use disorder, the most appropriate treatment is supportive care for AWS, including hydration, benzodiazepines, and possibly antipsychotics. Rapid correction of his sodium levels is not indicated due to the risk of osmotic myelinolysis. |
| Expertise Comment: ChatGPT incorrectly interpreted the patient’s symptoms as primarily related to Alcohol Withdrawal Syndrome (AWS) without adequately considering the patient’s laboratory results, particularly the severe hyponatremia (serum sodium 112 mEq/L). Although seizures can indeed accompany AWS, in this case, the diagnosis and treatment of hyponatremia should be the primary focus. Moreover, it incorrectly associated the risks of rapid correction of hyponatremia with cerebral edema. In fact, the main risk of rapidly correcting hyponatremia is osmotic myelinolysis, not cerebral edema. Osmotic myelinolysis results from rapid changes in blood osmolarity, leading to damage to the myelin sheath of brain neurons. | |
| Retrieve: A 48-year-old man is brought to the emergency department by his wife 20 minutes after she witnessed him vigorously shaking for about 1 minute. During this episode, he urinated on himself. He feels drowsy and has nausea. He has a history of chronic alcoholism; he has been drinking 15 beers daily for the past 3 days. Before this time, he drank 8 beers daily. His last drink was 2 hours ago. He appears lethargic. His vital signs are within normal limits. Physical and neurologic examinations show no other abnormalities. On mental status examination, he is confused and not oriented to time. Laboratory studies show: Hematocrit 44.0% Leukocyte count 12,000/mm3 Platelet count 320,000/mm3 Serum Na+ 112 mEq/L Cl- 75 mEq/L K+ 3.8 mEq/L HCO3- 13 mEq/L Urea nitrogen 6 mEq/L Creatinine 0.6 mg/dL Albumin 2.1 g/dL Glucose 80 mg/dL Urgent treatment for this patient’s current condition puts him at increased risk for which of the following adverse events?Ä: Cerebral edema B: Cardiac arrhythmia C: Hyperglycemia D: Osmotic myelinolysis E: Wernicke encephalopathy Correct Answer: D | |
| Q2: A 27-year-old gravida 2, para 1 comes… | |
| Answer: | |
ChatGPT: Based on the clinical presentation described, the most likely diagnosis for this patient is Rubella (German measles) infection during pregnancy, which can lead to congenital rubella syndrome (CRS) in the fetus if contracted during the first trimester. The typical symptoms of rubella include a maculopapular rash starting on the face and spreading downwards, along with lymphadenopathy and joint swelling. Given the risk of congenital rubella syndrome to the fetus, the most appropriate next step in management is: C) Vaccination Vaccination against rubella is contraindicated during pregnancy due to the risk of causing CRS. However, after delivery, the patient should receive the rubella vaccine to prevent future infections. This will not only protect the patient from future rubella infections but also prevent the risk of transmitting the virus to others, including vulnerable populations such as pregnant women. Options A, D, and E are not appropriate in this scenario. Oral acyclovir therapy (Option A) is used for herpes simplex virus infections, not rubella. Reassurance and symptomatic therapy (Option D) may provide comfort but do not address the potential risk to the fetus. Intravenous immunoglobulin therapy (Option E) is not indicated for rubella infection during pregnancy. Termination of pregnancy (Option B) is not necessary unless there are severe fetal abnormalities detected on further evaluation. |
Our: The patient’s history of a rash that started on her face and progressed downwards, along with the presence of postauricular lymphadenopathy, is suggestive of rubella infection. Rubella is a highly contagious viral infection that is spread through respiratory droplets or direct contact with infected secretions. It is characterized by a prodrome of malaise, low-grade fever, and lymphadenopathy, followed by the appearance of a pink maculopapular rash that begins on the face and spreads downwards. The rash is usually itchy and lasts for 3-5 days. Rubella infection during pregnancy, especially during the first trimester, can have serious consequences for the developing fetus. It can cause congenital rubella syndrome (CRS), which is characterized by a wide range of birth defects, including cataracts, deafness, heart defects, and intellectual disability. The most appropriate next step in management for this patient is termination of pregnancy. This is because the risk of CRS is high if the fetus is exposed to rubella during the first trimester, and the risk of birth defects outweighs the potential benefits of continuing the pregnancy. Options A, B, C, and E are not appropriate in this case. Acyclovir is used to treat herpes simplex virus infections, not rubella. Vaccination is not recommended during pregnancy due to the risk of fetal infection. Reassurance and symptomatic therapy are not sufficient to prevent the risk of CRS. Intravenous immunoglobulin therapy is not indicated for rubella infection during pregnancy. |
| Expertise Comment: ChatGPT made a made a basic error; it’s not appropriate to vaccinate during pregnancy. |
| 1 | Q3: A 1-year-old girl is brought to the physician… |
| Answer: | |
ChatGPT: Given the clinical presentation and findings described, this patient is most likely suffering from severe combined immunodeficiency (SCID), which is characterized by a deficiency in both T and B lymphocytes. The additional finding most likely to be present in this patient with SCID is: D) Decreased circulating T cells SCID is a primary immunodeficiency disorder characterized by profound defects in both cellular and humoral immunity. Patients with SCID typically have absent or markedly reduced T cell counts, as T cell development and function are severely impaired. This deficiency in T cells predisposes patients to recurrent, severe, and often life-threatening infections. While B cells may also be affected in SCID, the absence of T cells is the primary hallmark of the disease. Therefore, decreased circulating T cells (Option D) would be the most likely additional finding in this patient. Other options such as increased circulating IgE (Option B) or increased circulating neutrophils (Option F) are less specific to SCID and could be seen in a variety of other conditions. |
Our: Correct Answer: D Decreased circulating T cells This patient’s severe combined immunodeficiency (SCID) is most likely due to an autosomal recessive adenosine deaminase deficiency (the second most common cause of SCID), which results in disrupted purine metabolism and accumulation of metabolites (including deoxyadenosine and dATP) that are toxic to maturing lymphocytes. Accumulation of dATP also inhibits the function of ribonucleotide reductase, which is essential for the generation of deoxynucleotides, thereby impairing the proliferation of lymphocytes. These effects on lymphocytes proliferation and maturation lead to a decreased number of circulating T cells, B cells, and NK cells, resulting in both cellular and humoral immunodeficiency. Clinically, patients with SCID have no tonsils, no palpable lymph nodes, and a rudimentary thymus. In contrast, X-linked SCID (the most common variant) is caused by a defect in the IL-2 receptor, which is required for T-cell proliferation. Thus, X-linked SCID manifests with a decrease in circulating T cells and a normal number of B cells. Congenital immunodeficiency disorders Decreased levels of circulating parathyroid hormone in association with immunodeficiency can be seen in patients with DiGeorge syndrome. This syndrome also manifests with thymic hypoplasia and, possibly, the absence of tonsils. However, other features of CATCH-22 (e.g., congenital heart disease, dysmorphic facial features, hypocalcemia) would also be expected. Increased deoxyadenosine concentration in both the serum and urine and absent tonsils are not associated with DiGeorge syndrome. Increased circulating IgE levels can be seen in patients with Job syndrome, which is an immunodeficiency characterized by impaired neutrophil recruitment to the site of infection. Further manifestations would typically include coarse facial features, eosinophilia, severe eczema, and recurrent abscess formation. Increased deoxyadenosine concentration in both the serum and urine, absent tonsils, and an absent thymus shadow on x-ray are not associated with Job syndrome. Circulating B cells are decreased (not increased) in SCID due to adenosine deaminase deficiency and in X-linked agammaglobulinemia, an immunodeficiency characterized by a defective Bruton tyrosine kinase gene. Although X-linked agammaglobulinemia can also present with recurrent, severe pyogenic infections beginning > 4–6 months after birth, it is extremely rare in girls given its mode of inheritance. Absent thymic shadow, tonsillar hypoplasia, and elevated deoxyadenosine levels are also not consistent with this diagnosis. In X-linked SCID (the most common variant), the number of circulating B cells is normal. This patient’s severe combined immunodeficiency (SCID) is most likely due to an autosomal recessive adenosine deaminase deficiency (the second most common cause of SCID), which results in disrupted purine metabolism and accumulation of metabolites (including deoxyadenosine and dATP) that are toxic to maturing lymphocytes. Accumulation of dATP also inhibits the function of ribonucleotide reductase, which is essential for the generation of deoxynucleotides, thereby impairing the proliferation of lymphocytes. These effects on lymphocytes proliferation and maturation lead to a decreased number of circulating T cells, B cells, and NK cells, resulting in both cellular and humoral immunodeficiency. Clinically, patients with SCID have no tonsils, no palpable lymph nodes, and a rudimentary thymus. In contrast, X-linked SCID (the most common variant) is caused by a defect in the IL-2 receptor, which is required for T |
| Expertise Comment: ChatGPT’s explanation of the pathology is unclear, such as the specific type of SCID (Severe Combined Immunodeficiency caused by adenosine deaminase deficiency) and how this condition affects the proliferation and maturation of lymphocytes, leading to a reduced number of T cells. | |
| Retrieve: …The treatment for mild to moderate cases is mainly supportive, with early treatment of infections and prophylaxis with trimethoprim-sulfamethoxazole if needed, but severe cases require definitive therapy with hematopoietic stem cell transplantation. Congenital immunodeficiency disorders A defect in the interleukin-2 receptor subunit gamma gene (IL2RG) is the most common mutation associated with severe combined immunodeficiency (SCID), which can manifest soon after birth with recurrent episodes of otitis media, candidiasis, and sepsis. However, patients would also typically show failure to thrive and have a history of chronic diarrhea. A CBC, moreover, would reveal a low absolute lymphocyte count (< 1000/mm3). Delayed umbilical cord separation would not be expected. A defect in Bruton tyrosine kinase (BTK) is responsible for X-linked agammaglobulinemia (XLA), which may result in recurrent otitis media from infection with extracellular bacteria such as S. pneumoniae and H. influenzae. However, infants with B-cell defects such as XLA typically do not develop symptoms before 32̆0136 months because they generally retain passively acquired maternal antibodies up to this age. Delayed umbilical cord separation would also not be expected. Moreover, candidiasis is more commonly seen in defects involving T cells and granulocytes than in isolated primary defects of antibody production such as XLA. Defective NADPH oxidase results in chronic granulomatous disease (CGD), which may manifest in infancy with candidiasis (since C. albicans is catalase-positive) and neutrophilia during episodes of infection. However, these infections tend to remain localized, and sepsis is not as common as in other primary immune deficiency syndromes. Other pathognomonic features of CGD include GI obstruction and urinary retention caused by granulomas. Delayed umbilical cord separation would not be expected. Defective microtubules due to a loss of function of the lysosomal trafficking regulator gene (LYST) is the underlying pathophysiology of Chediak-Higashi syndrome (CHS), which can result in systemic candidiasis. However, patients with CHS would have neutropenia, not neutrophilia. Features such as oculocutaneous albinism and/or ocular motility disorders (e.g., strabismus, nystagmus) would also be present. Delayed umbilical cord separation would not be expected. A defectively organized actin cytoskeleton due to a mutation in the WAS gene is the underlying cause of Wiskott-Aldrich syndrome (WAS), which can manifest in infancy with recurrent otitis media as well as scaly erythematous lesions in the diaper area due to atopic dermatitis. However, the absence of thrombocytopenia on a CBC essentially rules out WAS. Delayed umbilical cord separation would also not be expected… |
| 1 | Q4: A 27-year-old woman presents to the office with concerns about her long struggle with her physical appearance since adolescence. She says she has always been l̈argeänd was constantly targeted by her classmates and coworkers for being so. Her main concern at the moment is her acne and unwanted facial hair on her upper lip, for which she often visits a local spa. She has tried numerous diet plans, exercise regimens, and cosmetic products with little to no effect. Recently, she underwent a glucose tolerance test that showed a plasma glucose level of 160 mg/dL (8.9 mmol/L) after 2 hours of a 75 g dose of oral glucose. She has a family history of type 2 diabetes mellitus and a menstrual cycle that occurs every 45 days. Her pulse is 72/min and the blood pressure is 138/80 mm Hg. On physical examination, her height is 160 cm (5 ft 3 in) and her weight is 85 kg (187 lb), and she has severe inflammatory acne over the cheeks and forehead and dark coarse hairs on the back. What is the most likely diagnosis of this patient? A: Hypothyroidism B: Idiopathic hirsutism C: Cushing’s disease D: Polycystic ovarian syndrome (PCOS) E: Ovarian hyperthecosis |
| Answer: | |
ChatGPT: The most likely diagnosis for this patient is: D: Polycystic ovarian syndrome (PCOS) PCOS is a common endocrine disorder in women of reproductive age, characterized by hyperandrogenism, irregular menstrual cycles, and polycystic ovaries on ultrasound. The clinical features seen in this patient, including acne, unwanted facial hair (hirsutism), obesity, irregular menstrual cycles, and a family history of type 2 diabetes mellitus, are all consistent with PCOS. Other options: A: Hypothyroidism - While hypothyroidism can present with weight gain and menstrual irregularities, it is less likely given the presence of hirsutism and obesity in this patient. B: Idiopathic hirsutism - This is unlikely as idiopathic hirsutism typically presents with hirsutism without other features of PCOS such as irregular menstrual cycles and obesity. C: Cushing’s disease - While Cushing’s disease can present with weight gain and hirsutism, it is less likely in this patient as there are no signs of hypercortisolism such as central obesity, buffalo hump, or purple striae. E: Ovarian hyperthecosis - This is a rare condition characterized by luteinization of ovarian stromal cells leading to hyperandrogenism. However, PCOS is a more common cause of hirsutism and irregular menstrual cycles in young women. |
Our: The patient’s history of being "large" since adolescence, a family history of type 2 diabetes mellitus, and a menstrual cycle that occurs every 45 days, along with her physical examination findings of severe inflammatory acne over the cheeks and forehead and dark coarse hairs on the back, are suggestive of polycystic ovary syndrome (PCOS). PCOS is a common endocrine disorder among women of reproductive age and is characterized by hyperandrogenism, chronic anovulation, and polycystic ovaries on ultrasound. The condition is associated with insulin resistance, which can lead to hyperglycemia and impaired glucose tolerance, as seen in this patient’s glucose tolerance test. The other options listed are not as likely given the patient’s presentation: A: Hypothyroidism is characterized by a slowing of metabolic processes, including growth and development, and is not associated with the symptoms or findings seen in this patient. B: Idiopathic hirsutism is a condition characterized by excessive hair growth in women, often on the face, chest, and back, but it does not typically present with the other symptoms or findings seen in this patient. C: Cushing’s disease is a condition caused by excessive production of cortisol, which can lead to weight gain, central obesity, and facial rounding, but it does not typically present with the other symptoms or findings seen in this patient. D: Polycystic ovarian syndrome (PCOS) is a common endocrine disorder among women of reproductive age and is characterized by hyperandrogenism, chronic anovulation, and polycystic ovaries on ultrasound. The condition is associated with insulin resistance, which can lead to hyperglycemia and impaired glucose tolerance, as seen in this patient’s glucose tolerance test. E: Ovarian hyperthecosis is a condition characterized by excessive production of androgens by the ovaries, which can lead to hirsutism and acne, but it does not typically present with the other symptoms or findings seen in this patient. |
| Expertise Comment: Our explanation is slightly more detailed in terms of linking the patient’s specific symptoms to the characteristics of PCOS and explaining why other diagnoses are less likely. It provides a more comprehensive understanding of PCOS and its relation to the patient’s symptoms and test results. On the other hand, ChatGPT offers a more concise but still accurate diagnosis. | |
Retrieve: The clinical features of PCOS are described here briefly but are reviewed in detail separately. (See "Clinical manifestations of polycystic ovary syndrome in adults".) PCOS is thought to be one of the most common endocrinopathies in women, affecting between 5 and 10 percent of women, depending upon the population studied [1]. (see "Epidemiology, phenotype, and genetics of the polycystic ovary syndrome in adults", section on ’Epidemiology’). The syndrome is characterized clinically by oligomenorrhea and hyperandrogenism, as well as the frequent presence of associated risk factors for cardiovascular disease, including obesity, glucose intolerance, dyslipidemia, fatty liver, and obstructive sleep apnea. Other features include:… (Robert L Barbieri and William F Crowley, 2024) |

### v3

@@
 Abstract
-With the explosive growth of medical data and the rapid development of artificial intelligence technology, precision medicine has emerged as a key to enhancing the quality and efficiency of healthcare services.
-In this context, Large Language Models (LLMs) play an increasingly vital role in medical knowledge acquisition and question-answering systems.
-To further improve the performance of these systems in the medical domain, we introduce an innovative method that jointly trains an Information Retrieval (IR) system and an LLM during the fine-tuning phase.
-This approach, which we call Joint Medical LLM and Retrieval Training (JMLR), is designed to overcome the challenges faced by traditional models in handling medical question-answering tasks.
-By employing a synchronized training mechanism, JMLR reduces the demand for computational resources and enhances the model’s ability to leverage medical knowledge for reasoning and answering questions.
-Our experimental results demonstrate that JMLR-13B (81.2% on Amboos, 61.3% on MedQA) outperforms models using conventional pre-training and fine-tuning Meditron-70B (76.4% on AMBOSS, 60.3% on MedQA).
-For models of the same 7B scale, JMLR-7B(68.7% on Amboos, 51.7% on MedQA) significantly outperforms other public models (Meditron-7B: 50.1%, 47.9%), proving its superiority in terms of cost (our training time: 37 hours, traditional method: 144 hours), efficiency, and effectiveness in medical question-answering tasks.
-Through this work, we provide a new and efficient knowledge enhancement tool for healthcare, demonstrating the great potential of integrating IR and LLM training in precision medical information retrieval and question-answering systems.
+Large Language Models (LLMs) have demonstrated a remarkable potential in medical knowledge acquisition and question-answering.
+However, LLMs can potentially hallucinate and yield factually incorrect outcomes, even with domain-specific pretraining.
+Previously, retrieval augmented generation (RAG) has limited success in addressing hallucinations.
+Unlike previous methods in RAG where the retrieval model was trained separately from the LLM, we introduce JMLR (for Jointly trains LLM and information Retrieval (IR)) during the fine-tuning phase.
+The synchronized training mechanism enhances JMLR’s ability to retrieve clinical guidelines and leverage medical knowledge to reason and answer questions and reduces the demand for computational resources.
+We evaluated JMLR on the important medical question answering application.
+Our experimental results demonstrate that JMLR-13B (70.5%) outperforms a previous state-of-the-art open-source model using conventional pre-training and fine-tuning Meditron-70B (68.9%) and Llama2-13B with RAG (54.9%) on a medical question-answering dataset.
+JMLR-13B (148 GPU hours) also trains much faster than Meditron-70B (42630 GPU hours).
+Through this work, we provide a new and efficient knowledge enhancement tool for healthcare, demonstrating the potential of integrating IR and LLM training for medical question-answering systems 111The code(https://github.com/believewhat/JMLR-Joint-Medical-LLM-and-Retrieval-Training), along with selected retrieval data that can be made public, is included in the supplementary material and will be made publicly accessible with CC-BY 4.0 license upon the paper’s acceptance..
 1 Introduction
-Informed clinical decision-making is fundamentally rooted in understanding and following the logical diagnosis chain.
-However, the expertise required to employ this diagnostic logic effectively is not universally accessible, posing a significant challenge in healthcare.
-Recent developments in Large Language Models (LLMs) offer potential solutions to democratize access to this essential aspect of medical knowledge (Achiam et al., 2023; Touvron et al., 2023; Singhal et al., 2022; Peng et al., 2023; Bommasani et al., 2021; Hoffmann et al., 2022; Raffel et al., 2020; Gao et al., 2020; Computer, 2023; Soldaini et al., 2023; Bubeck et al., 2023; Wei et al., 2022; Wang et al., 2022; Yang et al., 2023).
-LLMs’ historical development and evaluation predominantly focused on generalized tasks, with models pretrained from large amounts of data sourced from various online platforms.
-This approach, however, often resulted in a dearth of high-quality, domain-specific evidence, a shortfall particularly evident in the medical field Roziere et al.
+Effective clinical decision-making relies on a logical diagnostic chain, which requires specialized knowledge that isn’t widely accessible, presenting a major healthcare challenge.
+Large Language Models (LLMs) show promise in making this critical medical knowledge more accessible (Peng et al., 2023; Yang et al., 2023).
+Traditionally, LLMs have been developed for general tasks using data from diverse online sources, leading to a lack of high-quality, domain-specific information, especially in medicine Wornow et al.
+(2023); Singhal et al.
 (2022).
-A viable strategy to expand medical knowledge is to initiate training with medical-purpose LLMs and then refine this training with targeted domain-specific datasets.
-This hybrid training approach, blending broad linguistic capabilities with specific medical expertise, has been effectively applied in various medical models (Yunxiang et al., 2023; Zhang et al., 2023; Toma et al., 2023; Chen et al., 2023b).
+Previous researches expand medical knowledge by continue training general domain LLMs with domain-specific datasets.
+(Yunxiang et al., 2023; Zhang et al., 2023; Toma et al., 2023).
 However, LLMs face the significant challenge of “hallucination”, where models generate plausible but incorrect or unverified information (Ji et al., 2023; Bang et al., 2023).
-Such errors raise serious concerns in healthcare, where precision and reliability are imperative (Ahmad et al., 2023).
-Hence, recent works proposed retrieval augmented generation (RAG): they first train a retriever to obtain relevant documents from a large corpus based on the input query and then train an LLM to generate a response based on the input query and the retrieved documents (Lewis et al., 2020; Borgeaud et al., 2022a; Cheng et al., 2023; Xiong et al., 2024).
-By retrieving validated, up-to-date documents, RAG identifies relevant knowledge and provides contextual grounding for LLMs, thereby alleviating hallucination issues (Shuster et al., 2021; Zhu et al., 2023).
-Although RAG shows high accuracy in benchmarks for open-domain question answering, its effectiveness in specific domains is yet in doubt because retrievers trained in the general domain usually perform worse than those fine-tuned in the specific domain.
+Such errors raise serious concerns in healthcare, where accuracy is imperative (Ahmad et al., 2023).
+Hence, recent works proposed retrieval augmented generation (RAG): they first train a retriever to obtain relevant documents from a general domain corpus based on the input query and then train an LLM to generate a response based on the input query and the retrieved documents (Lewis et al., 2020; Borgeaud et al., 2022a; Cheng et al., 2023; Xiong et al., 2024).
+By retrieving domain specific documents, RAG identifies relevant knowledge and provides contextual grounding for LLMs, thereby alleviating hallucination issues (Shuster et al., 2021; Zhu et al., 2023).
+Although RAG shows high accuracy in benchmarks for open-domain question answering, its effectiveness in specific domains is yet in challenge because retrievers trained in the general domain usually perform worse than those fine-tuned in the specific domain (Gao et al., 2023; Zhao et al., 2022; Thakur et al., 2021).
 Fine-tuning retrievers requires document-query pairs, which may not be readily available for a specific domain.
@@
 Moreover, RAG trains retriever and LLM separately.
-When training LLM with a frozen retriever, it gathers documents it deems most relevant, which may not always align with what the LLM requires to produce the correct answer.
-To solve these challenges, we introduce Joint Medical LLM and Retrieval Training (JMLR), a novel approach that synchronizes the training of an LLM and retriever.
-We evaluate whether JMLR helps enhance reasoning and professional question-answering capability.
-Like RAG, JMLR enhances question answering by fetching relevant documents.
+When training LLM with a frozen retriever, the retrieval may not be optimized to align with LLM to produce the correct answer (Asai et al., 2024; Yoran et al., 2023; Rubin et al., 2022).
+In this study, we introduce a novel method to jointly train retrieval and LLM.
+Although the method is generalizable, we focus on the evaluation on the important medical domain where the accuracy of the generated content is crucial.
+Specifically, we introduce Joint Medical LLM and Retrieval Training (JMLR), a novel approach that jointly train LLM and retriever.
+As shown in Figure 2, JMLR presents a novel approach compared to the traditional pretrain-finetune process.
+JMLR enhances question answering by fetching relevant domain-specific documents that maximize the potential of LLMs.
 This involves including each retrieved document in the initial input question before the LLM’s response attempt, thereby augmenting the input to train the LLM for answer generation.
-In addition to RAG, JMLR introduces a unique mechanism called LLM-Rank loss to train the retriever.
+For training jointly, JMLR introduces a unique mechanism called LLM-Rank loss to train the retriever.
 This is achieved by evaluating the improvement in LLM’s performance upon including any candidate retrieved documents.
@@
 When reducing this loss, the retriever is trained to prioritize candidate documents that significantly aid the LLM.
-As shown in Figure 1, JMLR trains LLM and retriever jointly in the same step and thus improves medical question answering, especially in scenarios requiring nuanced understanding and specific information retrieval.
+Our results show that the joint training improves medical question answering, especially in scenarios requiring nuanced understanding and specific information retrieval.
 To validate the effectiveness of JMLR in terms of reducing computational resource requirements and fully utilizing given knowledge, we designed three experimental setups.
 These experiments evaluated whether direct fine-tuning, instead of traditional pretraining, would improve JMLR’s performance.
-We utilized external resources, including the MIMIC-IV dataset (Johnson et al., 2020), medical textbooks, and diverse medical guidelines for knowledge expansion.
-Our research presents a novel approach compared to the conventional pretrain-finetune process (initial pretraining followed by finetuning on a downstream task).
-Our technique has demonstrated superior results, achieving an impressive 53.8% accuracy on the MedQA dataset and 70.2% on AMBOSS, outperforming the traditional method’s 49.8% on MedQA and 52.8% on AMBOSS.
-Furthermore, our 13B model surpasses the performance of both the Meditron 70B and ChatGPT models in the USMLE and AMBOSS datasets’ question-answering tasks.
+We utilized external data resources for retrieval, including the MIMIC-IV dataset (Johnson et al., 2020), medical textbooks, and diverse medical documents for knowledge expansion.
+We experimented with different LLMs, including PMC-Llama, Meditron, ChatGPT with or without MedRAG.
+Our JMLR-7B has demonstrated superior results, achieving an impressive 62.45% accuracy, outperforming the traditional method’s 53.2% on the medical QA dataset.
+Furthermore, our 13B model (69.2%) surpasses the performance of both the open-source medical LLM (Meditron 70B: 68.9%) and closed-source general LLM (ChatGPT: 58.5%).
 Moreover, evaluation by both GPT-4 and domain experts supports the superior performance of JMLR.
 We summarize our key contributions as follows:
-We propose a new method 111The code, along with selected retrieval data that can be made public, is included in the supplementary material and will be made publicly accessible upon the paper’s acceptance.
-that integrates retrieval and LLM training together, resulting in JMLR surpassing the state-of-the-art open source model by 12%.
-Our training method decreases the training time required for knowledge expansion to just 37 hours.
-This is a considerable reduction compared to the conventional pretrain-finetune process, where pretraining alone takes 127 hours and finetuning adds an additional 17 hours.
-JMLR, compared to the public SOTA models, is capable of providing enhanced reasoning for improved explainability, and the performance is better than ChatGPT(winning rate is 0.63).
-2 Related Work
-Medical Large Language Model
-Medical LLMs have evolved from adapting existing models like BERT using biomedical datasets (Gu et al., 2021; Lee et al., 2020), to integrating document linking (Yasunaga et al., 2022b) and knowledge graphs (Yasunaga et al., 2022a).
-Decoder-only architectures such as GPT and Llama were then trained with domain-specific data (Radford et al., 2018; Touvron et al., 2023; Soldaini et al., 2023; Wu et al., 2023; Gema et al., 2023; Yunxiang et al., 2023; Zhang et al., 2023; Labrak et al., 2024) or some synthetic data (Tran et al., 2023; Han et al., 2023; Kweon et al., 2023; Wang et al., 2023b).
-Recently, scaling up data and model parameters led to models like GatorTronGPT (Peng et al., 2023) and Clinical-Camel (Toma et al., 2023), with studies on Flan-PaLM and PaLM-2’s medical reasoning (Singhal et al., 2023a, b).
-Furthermore, researchers have developed MEDITRON (Chen et al., 2023b), an open-source suite of medical LLMs with 7B and 70B parameters, offering enhanced performance compared to state-of-the-art baselines and outperforming closed-source models like GPT-3.5 and Med-PaLM, while closely rivaling GPT-4 and Med-PaLM-2 in effectiveness.
-Nonetheless, the persistent augmentation in the model’s size has not yielded commensurate enhancements in its performance metrics.
-This parameter expansion fails to adequately remedy the phenomenon of model forgetting (Wang et al., 2023a; Luo et al., 2023), and it appears that the model’s parameters are insufficient for retaining extensive knowledge.
-In contrast to existing approaches, we combine retriever and LLM in our training process to mitigate the forgetting issue.
-Our work provides two versions of medical LLMs: 7B and 13B.
-Evaluation results indicate that our models surpass previous traditional pre-trained language models in performance, and our 13B model even outperforms the MEDITRON 70B model.
-Retrieval-Augmented Language Models
-In the clinical domain, the occurrence of factual inaccuracies by language models (LMs), attributed to the gap in medical knowledge, presents considerable risk, with potentially severe implications such as misdiagnoses Petroni et al.
-(2019); Sung et al.
-(2021); Yao et al.
-(2022); Singhal et al.
-(2023a).
-Enhancing LMs by incorporating pertinent information sourced from diverse knowledge repositories has been demonstrated as an effective method for boosting performance across multiple Natural Language Processing tasks, notably in the realm of language modeling (Min et al., 2022).
-The retrieval component remains static in traditional RAG (Retrieval-Augmented Generation) models, while updates are applied solely to the LLM.
-Specifically, the input is utilized as a query in a two-step process.
-Initially, a retrieval system selects a collection of documents, which are essentially sequences of tokens, from a designated corpus.
-Subsequently, a language model assimilates these retrieved documents, employing them as supplementary information to formulate a conclusive prediction (Yu, 2022; Izacard et al., 2022).
-However, merely integrating the documents retrieved into the LM does not necessarily mean that the LM can use external knowledge.
-Consequently, several researchers have shifted the focus towards devising methods for guiding models in applying external knowledge.
-For instance, Atlas (Izacard et al., 2022) enhances an encoder-decoder model in tandem with the retriever by treating documents as latent variables.
-RETRO (Borgeaud et al., 2022b) modifies the decoder-only architecture to include retrieved texts and undertakes the pretraining of the LM from the ground up.
-However, LLMs are not inherently structured to seamlessly integrate retrieved content, often leading to discrepancies or inconsistencies when blending pre-trained model responses with external information.
-Fusing externally retrieved information with the model’s internal knowledge base can be intricate and flawed.
-This sometimes results in an awkward or inconsistent amalgamation of information, where the LLM may struggle to incorporate the external data into its responses coherently.
-Even though some researchers have considered training retriever on unsupervised data so that it can provide documents that are well-matched with LLM, such as REALM (Guu et al., 2020), REPLUG (Shi et al., 2023) and RA-DIT (Lin et al., 2023), these retriever systems are trained independently rather than in synchronization with LLM training.
-Our method integrates and synchronously trains the retriever and LLM, enabling the retriever to return documents more suited to LLM, thus enhancing LLM’s performance.
-Additionally, we have also conducted comparisons with step-by-step training methods to prove the effectiveness of our method.
-3 Materials and Methods
-There is a notable shortfall in biomedical knowledge datasets designed for training instruction-compliant assistants.
-To address this, we have collected the dataset of biomedical knowledge resources, primarily sourced from a wide array of clinical guidelines and textbook data, for pretraining.
-For the fine-tuning phase, we have gathered an extensive collection of medical QA datasets, aimed at augmenting the question-answering capabilities of LLMs.
-3.1 Datasets
-Knowledge Sources:
-Data quality is important for retrieval of JMLR and for pretraining and finetuning LM.
-In routine practice, clinicians typically encounter at least one patient-care-related question for every two patients they see.
-They often turn to guidelines like UpToDate 222www.uptodate.com.
-UpToDate is an expert-authored, evidence-based clinical knowledge system that helps clinicians to make point-of-care decisions.
-Other high-quality resources include textbooks and electronic health record (EHR) notes.
-Our collection of UpToDate, textbooks and MIMIC-IV notes (the only publicly available EHR notes) has been used as documents to retrieve.
-The fine-tuning process is based on the pre-trained Llama model, where we leveraged datasets from two significant sources: the USMLE and AMBOSS, which are essential for preparing for the USMLE examinations.
-The USMLE dataset, sourced directly from the official USMLE website, included a range of sample questions for Step1, Step2CK, and Step3, released between June 2022 and March 2023.
-These questions represent the complex medical knowledge and ethical scenarios that medical students and practitioners are expected to navigate.
-Additionally, we utilized the AMBOSS question bank, a comprehensive resource widely used by medical professionals and students.
-This dataset provided an extensive array of Step1, Step2CK, and Step3-type questions, further enriching our training material with practical and diverse medical scenarios.
-The MedMCQA dataset comprises 194k multiple-choice questions from Indian medical entrance exams, covering 2.4k healthcare topics and 21 medical subjects (Pal et al., 2022).
-The MMLU dataset Hendrycks et al.
-(2020), covering 57 subjects, including STEM and social sciences, features four-option multiple-choice questions for each subject.
-We focused on nine subjects most pertinent to medical and clinical knowledge — high school biology, college biology, college medicine, professional medicine, medical genetics, virology, clinical knowledge, nutrition, and anatomy.
-3.2 Joint Medical LLM and Retrieval Training (JMLR)
-3.2.1 LLM
-The original Llama model supports a maximum of 4k tokens, which poses a challenge when multiple documents are retrieved.
-To address this, we utilized the Shifted Sparse Attention (S2-Attn) mechanism (Chen et al., 2023a), which can be used to process long sequences by mitigating the high memory cost and slow processing time associated with standard self-attention in LLMs.
-S2-Attn achieves this by dividing the input into several sequence spans within the self-attention modules and then introducing a shifted pattern to facilitate information exchange between these spans.
-This approach efficiently extends the context length without incurring additional computational costs and demonstrates performance close to the standard attention baseline.
-The method, designed for efficient training, retains the standard self-attention architecture during inference, making it compatible with existing infrastructure and optimizations.
-3.2.2 Retriever
+We propose a new method that jointly train retrieval and LLM, resulting in JMLR surpassing the state-of-the-art open source model.
+By training jointly, we do not require additional human annotation, thereby enhancing flexibility in domain or downstream task adaptation.
+When training 7B parameter model with a single Nvidia A100 GPU, our training method decreases the training time required for knowledge expansion to just 100 hours.
+This is a considerable reduction compared to Meditron’s pretrain-finetune process, where pretraining alone takes 588 hours, and finetuning adds an additional 36 hours.
+2 Problem Formulation & Traditional Method
+Given a set of medical questions and a set of medical documents containing medical knowledge, our goal is to construct a language model that can provide accurate answers.
+Formally, for each question , there is a corresponding correct answer within a set of options , where . The model maps each question to a predicted answer:
+where represents the parameters of the model.
+Our objective is to find the optimal parameters that minimize the loss function , which measures the discrepancy between the predicted answer and the correct answer :
+The loss function can be instantiated as a cross-entropy loss for classification tasks, where the number of classes is equal to the number of options.
+To solve this problem, previous methods pretrained LLM to learn medical knowledge for a medical LLM, and then finetuned medical LLM on medical QA task (Chen et al., 2023b; Wu et al., 2023; Yunxiang et al., 2023; Toma et al., 2023).
+Specifically, such traditional pretrain-finetune pipeline first continued pretraining general domain LLMs on medical documents with next-token prediction loss function, and then finetuned the medical LLMs to select given question and option description with loss function .
+3 Joint Medical LLM and Retrieval Training (JMLR)
+In comparison to traditional method, JMLR did not pretrain LLM on medical documents.
+Instead, JMLR selected question-related medical documents as additional context during fine-tuning of an open-domain LLM.
+Specifically, given a question , we trained retriever to find helpful medical documents from , which was then used to train LLM to generate the final answer .
+3.1 Retriever
 For information retrieval we used the ColBERT model and the trained weight as the initial model.
@@
 The resulting embeddings for queries () and documents () are computed as follows: is normalized after being processed through a CNN layer applied to BERT’s output, and undergoes a similar process but includes a filtering step to remove certain embeddings.
-3.2.3 JMLR Architecture
+3.2 LLM
+We adapted Llama as our LLM model.
+The original Llama model supports only a maximum of 4k tokens, which poses a challenge when multiple documents are retrieved.
+To address this, we utilized the Shifted Sparse Attention (S2-Attn) mechanism (Chen et al., 2023a), which can be used to process long sequences by mitigating the high memory cost and slow processing time associated with standard self-attention in LLMs.
+S2-Attn divides input into sequence spans within self-attention modules and introduces shifted patterns for inter-span information exchange.
+This extends context length efficiently without extra computational costs and achieves near-baseline performance.
+It maintains the standard self-attention architecture during inference, ensuring compatibility with existing infrastructure and optimizations.
+3.3 JMLR Architecture
 However, Colbert matches queries and documents based on the retriever similarity, returning scores and documents.
 To enable the retriever system to return documents that are more helpful for providing accurate answers to Llama, we constructed the JMLR architecture.
-This structure employs a rank loss, , which links the Llama and retriever systems.
-The updating of retriever parameters is influenced by Llama’s content quality.
+Let be a set of queries, be a corpus of domain specific documents, and be the set of options.
+Our task is to construct a function parameterized by , which maps a question and a set of documents to an answer .
+The objective is to simultaneously learn the optimal parameters and that minimize the combined loss on the retrieved documents and the generated answers, formalized as: where is a loss function that evaluates the correctness of the answer and the relevance of the retrieved documents.
+The retrieval function is defined by:
+and the similarity function scores how relevant each document is to the question .
+The combined optimization involves updating both and through gradient descent to minimize the expected loss:
+where denotes the gradient of the loss function with respect to both sets of parameters.
+This dual-parameter optimization ensures that the retrieval function is aligned with the needs of the answer function, enhancing both the accuracy of the responses and the relevance of the information retrieved.
+To solve this, JMLR structure employs a rank loss, , which links the Llama and retriever systems.
+Llama’s content quality influences the updating of retriever parameters.
 If the returned document reduces Llama’s loss function more than document , it indicates that document is more helpful for answering than document .
@@
 in .
-For JMLR model training, the dataset comprises QA pairs from AMBOSS and USMLE, with the input being a question and the output being the correct choice, including the reasoning part.
+where is calculated as:
+and represents the contrastive logits:
+with the associated decision-based loss differences:
+For JMLR model training, the dataset comprises QA pairs, with the input being a question and the output being the correct choice, including the reasoning part.
 The top 7 documents with the highest scores are aggregated and fed into the model during the training process.
@@
 During each iteration, 30 scores are generated by the retriever, and we perform weighted random sampling based on these scores, with higher scores having a greater probability of selection.
-This approach ensures that each document inputted into the LLM (Large Language Model) is dynamic.
-The LLM will likely encounter different useful guidelines and some that are not useful.
+This approach ensures that each document inputted into the LLM model is dynamic.
+The LLM will likely encounter different useful documents and some that are not useful.
 This step enhances the LLM’s robustness, helping it discern which knowledge from which documents is useful and which documents are irrelevant.
+4 Experiment
+4.1 Dataset
+Medical document:
+Data quality of medical documents is important for LLMs to answer medical questions.
+Previous work primarily sourced from a wide array of medical research articles and clinical guidelines.
+Research papers, like those found in PubMed, provide foundational and current information on healthcare.
+Clinical guidelines are thoroughly designed protocols developed to assist clinicians in making decisions given patient medical history.
+Other high-quality medical documents to retrieve include medical textbooks.
+| Corpus | #Doc. |
+| PubMed | 23.9M |
+| Textbooks | 18 |
+| Cancer Care Ontario | 87 |
+| Center for Disease Control and Prevention | 621 |
+| Canadian Medical Association | 431 |
+| International Committee of the Red Cross | 49 |
+| National Institute for Health and Care Excellence | 1.7k |
+| Strategy for Patient-Oriented Research | 217 |
+| World Health Organization | 223 |
+| WikiDoc | 33k |
+Medical QA:
+The MedQA dataset Jin et al.
+(2021), sourced directly from the official USMLE website, included a range of sample questions for Step1, Step2CK, and Step3, released between June 2022 and March 2023.
+These questions represent the complex medical knowledge and ethical scenarios that medical students and practitioners are expected to navigate.
+Additionally, Amboss question bank is a comprehensive resource widely used by medical professionals and students.
+This dataset provided an extensive array of Step1, Step2CK, and Step3-type questions, further enriching our training material with practical and diverse medical scenarios.
+The MedMCQA dataset comprises 194k multiple-choice questions from Indian medical entrance exams, covering 2.4k healthcare topics and 21 medical subjects (Pal et al., 2022).
+MMLU-Medical is selected from the MMLU dataset Hendrycks et al.
+(2020), focused on nine subjects most pertinent to medical and clinical knowledge — high school biology, college biology, college medicine, professional medicine, medical genetics, virology, clinical knowledge, nutrition, and anatomy.
+Since this dataset has no training set, we opt to evaluate it using the LLM originally trained on MedMCQA.
+We individually apply supervised finetuning on each QA dataset and subsequently assess their performance on the corresponding test sets, unless otherwise specified.
+MedQA and Amboss both offer detailed reasoning traces, such as extensive answers or explanations.
+4.2 Baselines
+For comparison with traditional domain pretraining method, we select several public available medical domain pretrained LLMs as baselines: (1) PMC-Llama, a Llama model further pretrained on medical research articles from PubMed Central (Wu et al., 2023); (2) Meditron, a LLM pretrained on clinical guidelines and research papers (Chen et al., 2023b).
+To validate the benefit of joint training, we also compare JMLR with its naive version: RAG, which freezes the retriever and fetches the same document given a specific question.
 4.3 Training Details
@@
 We implement a cosine learning rate schedule, incorporating a warmup phase that accounts for 10% of the training duration and decays the learning rate to 10% of its peak value.
-In alignment with the practices outlined in Llama 2-chat touvron2023b, our training employs a learning rate of , a weight decay factor of 0.1, and manages a batch size of 2.
+In alignment with the practices outlined in Llama 2-chat Touvron et al.
+(2023), our training employs a learning rate of , a weight decay factor of 0.1, and manages a batch size of 2.
 The finetuning phase spans 5 epochs for all iterations.
@@
 Throughout these experiments, we utilize four A100 80 GB GPUs.
-The total training duration for our method, when applied to the MedQA and AMBOSS datasets, amounts to 37 hours.
-Additionally, the pretraining phase on medical guidelines takes approximately 127 hours, while the finetuning phase is approximately 17 hours on 7B Llama.
+The total training duration for our method amounts to 25 hours.
+Additionally, the pretraining phase on medical guidelines takes approximately 138 hours, while the finetuning phase is approximately 17 hours on 7B Llama.
 This comprehensive training approach ensures the effective adaptation and optimization of the model for specific medical applications.
 5 Main Results
-4.1 Effective of Retrieval
-Setup: For benchmarks with available public training sets, we individually apply supervised finetuning using JMLR on each set and subsequently assess their performance on the corresponding test sets.
-PubMedQA and MedMCQA both offer detailed reasoning traces, such as extensive answers or explanations.
-For MedQA, AMBOSS, and MedMCQA, we employ an alternative training set equipped with human-crafted explanations for each query.
-In this scenario, JMLR is trained to attach the answer directly to the explanation.
-Regarding MMLU-Medical, a dataset without its own training set, we opt to evaluate it using the JMLR model originally trained on MedQA and AMBOSS.
-We calculate the accuracy of different medical QA datasets based on different models.
-| Datasets | MMLU-Medical | MedMcQA | MedQA | AMBOSS |
-| PMC-Llama-7B | 59.7 | 57.6 | 42.4 | 43.7 |
-| Llama 2-7B | 56.3 | 54.4 | 44.0 | 46.5 |
-| Meditron-7B | 55.6 | 59.2 | 47.9 | 50.1 |
-| JMLR-7B | 57.2 | 61.3 | 51.7 | 68.7 |
-Results: In Table 1, our JMLR-7B model, operating at the 7B scale, demonstrates its superiority by outperforming a variety of state-of-the-art (SOTA) public models across several benchmark datasets in the medical field.
-The datasets include MMLU-Medical, MedMcQA, MedQA, and AMBOSS.
+| Model | Parameter | Open | MedQA | Amboss | MMLU | MedMCQA | Average |
+| Access | |||||||
+| GPT-4 | ? | No | 74.7 | 82.1 | 88.4 | 69.5 | 78.6 |
+| ChatGPT | ? | No | 50.2 | 49.1 | 69.4 | 51.0 | 54.9 |
+| Meditron | 70B | Yes | 60.7 | 76.4 | 73.6 | 65.1 | 68.9 |
+| RAG | 13B | Yes | 59.9 | 76.9 | 69.9 | 64.2 | 67.7 |
+| JLMR | 13B | Yes | 62.5 | 81.2 | 72.8 | 65.5 | 70.5 |
+As shown in Table 2, our JLMR with 13B parameters outperforms previous SOTA open access model Meditron 70B parameters and even closed access model such as ChatGPT.
+In the following sections, we will explain with more detailed analysis.
+5.1 Domain Retrieval > Domain Pretrain
+| Model | MedQA | Amboss | MMLU | MedMCQA | Average |
+| Llama 2 | 44.0 | 46.5 | 56.3 | 54.4 | 50.3 |
+| PMC-Llama* | 42.4 | 43.7 | 59.7 | 57.6 | 50.8 |
+| Meditron* | 47.9 | 50.1 | 55.6 | 59.2 | 53.2 |
+| RAG# | 54.6 | 70.7 | 62.1 | 62.4 | 62.5 |
+| JLMR# | 56.2 | 68.3 | 64.3 | 62.4 | 62.8 |
+Domain retrieval models (both RAG-7B and our JMLR-7B) outperform domain pretrained LLMs across all medical QA benchmarks, as shown in Table 3.
+In the MedQA and Amboss datasets, where JMLR-7B not only outperforms baselines but does so with a notable margin, scoring 56.2% in MedQA and 71.2% in Amboss.
+This is particularly significant when compared to Meditron-7B, which scores 47.9% and 50.1% in these datasets, respectively.
+The trend continues in MMLU-Medical and MedMcQA datasets, JMLR-7B achieved scores of 64.3% and 62.4%, respectively, surpassing the scores of PMC-Llama-7B (59.7% and 57.6%), and Meditron-7B (55.6% and 59.2%).
+Overall, the results illustrate that JMLR-7B model, on average, achieves about a 14% improvement in performance over its closest competitor, Meditron-7B.
+This highlights the effectiveness of our training method and the substantial advancements it brings to the field of medical benchmarking.
 The enhanced performance of JMLR-7B can be partially attributed to its use of Llama-2 as the foundational model, showcasing significantly higher average performance than other pre-trained baselines.
-However, the distinct edge comes from our integrated training approach that combines retriever and language model learning (LLM).
-This methodology not only further boosts the performance of Llama-2 in medical benchmarks but also enables the retriever component to effectively assist the LLM in sourcing relevant and helpful documents for answering questions.
-This synergy significantly reduces instances of hallucination and enhances the overall effectiveness of the model.
-In the MMLU-Medical and MedMcQA datasets, JMLR-7B achieved scores of 57.2% and 61.3%, respectively, surpassing the scores of PMC-Llama-7B (59.7% and 57.6%), Llama 2-7B (56.3% and 54.4%), and Meditron-7B (55.6% and 59.2%).
-The trend continues in the MedQA and AMBOSS datasets, where JMLR-7B not only outperforms these models but does so with a notable margin, scoring 51.7% in MedQA and an impressive 68.7% in AMBOSS.
-This is particularly significant when compared to Meditron-7B, which scores 47.9% and 50.1% in these datasets, respectively.
-Overall, the results illustrate that the JMLR-7B model, on average, achieves about a 10% improvement in performance over its closest competitor, Meditron-7B.
-This highlights the effectiveness of our training method and the substantial advancements it brings to the field of medical benchmarking.
-4.2 Effective of Joint LLM and Retrieval Training
-Setup: In our experiment, we compared the 7B and 13B Llama models.
-We employed conventional RAG, fine-tuning the 7B and 13B models(RAG-7B and RAG-13B).
-During training, we didn’t update the parameters of the retriever; only the parameters of Llama were updated.
-We also separated the training steps: initially, we fixed the parameters of the retriever and only updated those of Llama, and then fixed Llama’s parameters while updating the retriever’s using rank loss, resulting in JLMR separate models, JLMR-separate-7B and JLMR-separate-13B.
-Finally, we trained JLMR using our integrated approach to obtain JMLR-7B and JMLR-13B.
-Additionally, we compared these models with larger-scale models such as ChatGPT and Meditron-70B.
-Results: As shown in Figure 2, while the traditional RAG method offers some model accuracy enhancement, its impact is not particularly substantial.
-However, when we update the retriever using our rank loss(our-separate-7B and our-separate-13B), there is a noticeable improvement over the traditional RAG.
-Our-separate-7B and our-separate-13B achieved 51.8% and 65.4% in MedQA, and 54.5% and 71.3% in AMBOSS, respectively, outperforming the RAG variants.
-This improvement stems from the retriever’s enhanced ability, after the update, to identify documents that aid Llama in accurately answering questions.
-However, asynchronous training can be a bottleneck, especially in the efficient use of medical guidelines.
-While updating Llama’s parameters, the static nature of the retriever’s guidelines means that Llama might miss out on some crucial guidelines.
-This gap leads to the underutilization of potentially beneficial medical guidelines.
-Our joint training process,(JMLR-7B and JMLR-13B), effectively mitigates this issue, leading to a significant boost in performance.
-Because the parameters of the retriever are continuously updated during our training method, the best documents returned also change, allowing the model to encounter different documents that are helpful in answering questions, thereby enabling it to make better use of these medical guidelines.
-Compared to the larger models, JMLR-13B not only outperforms ChatGPT (which scores 55.4% in MedQA and 55.6% in AMBOSS) but also surpasses Meditron-70B.
-Specifically, JMLR-13B achieves an impressive 61.3% in MedQA and 81.2% in AMBOSS, indicating a substantial leap in model effectiveness and underscoring the success of our joint training approach.
-4.3 Effective of Diverse Training Processes
-Setup: To explore the impact of pretraining, finetuning, and information retrieval on model performance, we designed the following experiments: We tested the original Llama 2-7B on MedQA and AMBOSS datasets (Llama-7B).
-We also added a retriever during the inference stage to extract documents to assist Llama 2-7B in answering questions(Llama-7B+Retriever).
-Additionally, we finetuned Llama 2-7B on different training sets to obtain FT-Llama-7B.
-The same operation was repeated, adding a retriever during the inference stage to extract relevant documents to assist the model in answering questions(FT-Llama-7B+Retriever).
-Furthermore, to investigate the effectiveness of pretraining, we pretrained Llama 2 on medical guidelines(Llama-7B-Pretrain) and repeated the tests with and without adding a retriever (FT-Llama-7B-Pretrain, FT-Llama-7B-Pretrain+Retriever).
-Finally, we applied our training approach for further training on Llama-7B and Llama-7B-Pretrain (JMLR-7B, JMLR-7B-Pretrain).
-Result: Figure 3 displays the varied performance of models at different stages.
-We observed that Llama-7B+Retriever, compared to Llama 2-7B, did not improve but decreased effectiveness.
-This suggests that Llama 2-7B does not utilize these guidelines effectively, and longer guidelines may even lead to misunderstandings and hallucinations.
-In contrast, FT-Llama-7B significantly improved over Llama-7B, indicating that the finetuning phase effectively enhanced Llama’s medical QA capabilities.
-However, this phase did not aid in the utilization of medical guidelines.
-FT-Llama-7B-Pretrain further improved upon FT-Llama-7B, and adding Retriever during the inference stage resulted in a slight increase in performance.
-This improvement is likely due to the pretraining phase expanding Llama 2-7B’s medical knowledge, thereby enabling it to utilize medical guidelines more effectively.
-Lastly, JMLR-7B and JMLR-7B-Pretrain achieved the best results.
-However, the minor difference between JMLR-7B and JMLR-7B-Pretrain suggests that our method has already sufficiently expanded Llama 2-7B’s medical knowledge.
-Considering training time, JMLR-7B required only 37 hours, while JMLR-7B-Pretrain needed 127 hours (for pretraining) plus 37 hours (for our method), indicating that our approach is more efficient in expanding the model’s medical knowledge.
-4.4 Effective of the reasoning ability
-Setup: To verify that our model can accurately answer questions and provide detailed reasoning, we tasked JMLR-13B with generating reasons for its responses to the AMBOSS and MedQA test sets.
+However, the distinct edge comes from our integrated training approach that combines retriever and LLM.
+This methodology not only further enhances Llama-2’s performance in medical benchmarks but also equips the retriever component with the ability to effectively source relevant and beneficial documents to aid the LLM in answering questions.
+In contrast, ChatGPT, when devoid of medical guidelines, has been known to make basic mistakes, such as recommending vaccinations to pregnant women—a practice that contradicts medical norms as documented in Table 5.
+Our model, with the support of guidelines, avoids such errors.
+This synergistic approach substantially reduces hallucinations and bolsters the model’s overall reliability.
+5.2 Joint Training > RAG
+| Llama-7B | Llama-7B-ColBERT | FT-Llama-7B | FT-Llama-7B-ColBERT | JMLR-7B-Seperate | JMLR-7B | |
+| MedQA | 24.98 | 24.01 | 44.00 | 40.60 | 55.30 | 56.20 |
+| Amboss | 17.10 | 35.90 | 46.50 | 45.60 | 71.00 | 71.20 |
+During fine-tuning, synchronously updating both the Retriever and LLM yields better results than updating only the LLM.
+Table 2 shows that JLMR models significantly outperform traditional RAG methods.
+For example, JLMR-13B’s scores (72.8% on MMLU, 65.5% on MedMcQA) is far superior to RAG-13B (69.9% on MMLU, 64.2% on MedMcQA).
+This result suggests that without updating the retriever, the documents retrieved may be relevant to the question but not necessarily helpful to the LLM’s response.
+Joint training ensures that the retriever learns which documents are beneficial for Llama-2.
+To evaluate the adaptability of the JLMR retriever, we further applied it to other LLMs.
+Specifically, we employed different retrievers to identify relevant documents, which were then provided to ChatGPT for generating the final answer.
+The retriever trained with JMLR attained an average accuracy of 57.5%, while the retriever from RAG achieved an accuracy of 56.7%.
+These results indicate that the JLMR retriever effectively identifies documents that not only support Llama but also aid ChatGPT in selecting the correct answers.
+Interestingly, this experimental outcome aligns with recent studies on weak-to-strong generalization Burns et al.
+(2023), such as the Aligner Ji et al.
+(2024).
+It appears that training a retrieval module using a weaker model through the JMLR process enables it to function as a model-agnostic plug-and-play module, which subsequently enhances the performance of a stronger model.
+This suggests a promising research direction for future studies.
+6 Ablation Study
+To assess the effects of fine-tuning and information retrieval on model performance, we structured our experiments as follows:
+Original Model Evaluation: Test the original Llama 2-7B on the MedQA and Amboss datasets.
+Retrieval Augmentation: Evaluate the addition of a retriever during the inference stage to assist Llama 2-7B (Llama-7B+Colbert) in answering questions, analyzing the impact of the Colbert retriever.
+Fine-Tuning: Fine-tune Llama 2-7B on the Amboss and MedQA training sets to develop FT-Llama-7B.
+Retrieval with Fine-Tuning: Add a retriever during the inference stage to assist the fine-tuned model (FT-Llama-7B+ColBERT), to explore the effects of Colbert on the fine-tuned model.
+Asynchronous Training: Fix the parameters of the retriever while only updating those of Llama, then reverse to update only the retriever’s parameters using rank loss, resulting in JMLR-7B-Separate.
+Joint Training: Train the retriever and Llama simultaneously (JMLR-7B) to compare against separate training and to examine the impact of joint training on enhancing model performance.
+Table 4 displays the varied performance metrics of the models at different experimental stages. Observations are as follows:
+Llama-7B+Retriever (24.01% on MedQA, 35.90% on Amboss) underperformed compared to the original Llama 2-7B (24.98% on MedQA, 17.10% on Amboss), suggesting that the original Llama 2-7B model struggles to effectively utilize external documents, with longer documents potentially leading to misunderstandings or erroneous information synthesis.
+The FT-Llama-7B (44.0% on MedQA, 46.5% on Amboss) showed significant improvements over Llama-7B, indicating that the fine-tuning phase notably enhanced Llama’s capabilities in medical QA.
+However, this phase did not effectively leverage medical guidelines.
+The JMLR-7B-Separate model achieved superior performance (55.3% on MedQA and 71.0% on Amboss), surpassing the FT-Llama-7B-ColBERT (40.6% on MedQA, 46.5% on Amboss).
+The distinctive fine-tuning approach of JMLR-7B-Separate, which integrates medical documents during the tuning phase, enables the model to utilize medical resources more effectively.
+Finally, JMLR-7B (56.2% on MedQA, 71.2% on Amboss) uses a joint training approach that facilitates the retrieval of highly relevant medical documents, enhancing the model’s ability to accurately answer questions by presenting similar case-based treatment plans for consideration.
+As illustrated in Table 5, JMLR retrieves closely related cases, significantly improving the model’s decision-making process by offering comparable treatment options.
+7 Discussion: JMLR Effectiveness of the reasoning ability
+To verify that our model can accurately answer questions and provide detailed reasoning, we tasked JMLR-13B with generating reasons for its responses to the Amboss and MedQA test sets.
 Similarly, we had ChatGPT generate corresponding reasons.
-We employed two methods for scoring these reasons: 1.
-We used GPT-4 to judge which model’s generated reasoning was better based on the ground truth.
-2.
-We engaged three medical professionals to review and assess a sample of 20 questions drawn from the test sets.
+We employed two methods for scoring these reasons: 1) We used GPT-4 to judge which model’s generated reasoning was better based on the ground truth.
+2) We engaged three medical professionals to review and assess a sample of 20 questions drawn from the test sets.
 These doctors were asked to judge which model’s reasoning was closer to the actual, clinically accurate reasons.
-Result: Table 2 demonstrates the varying win rates of our model under two different evaluation methods.
 In the assessment using GPT-4, JMLR-13B achieved a win rate of 0.63, while under the evaluation of medical professionals, our model attained a win rate of 0.60.
@@
 The close alignment of the win rates from both GPT-4 and medical professionals’ evaluations further underscores the model’s robustness and reliability in medical question-answering and reasoning tasks.
-| Comparison | Win Rate |
-| GPT4 | 0.63 |
-| Expertise | 0.6 |
+8 Related Work
+Medical Large Language Model
+Medical LLMs have advanced from adapting models like BERT with biomedical datasets (Gu et al., 2021; Lee et al., 2020) to incorporating knowledge graphs (Yasunaga et al., 2022).
+Subsequently, architectures such as GPT (Bubeck et al., 2023) and Llama (Touvron et al., 2023) were trained on domain-specific (Wu et al., 2023; Gema et al., 2023; Yunxiang et al., 2023; Zhang et al., 2023; Labrak et al., 2024) or synthetic data (Tran et al., 2023; Han et al., 2023; Kweon et al., 2023).
+Recent efforts have scaled up data and model parameters, resulting in GatorTronGPT (Peng et al., 2023) and Clinical-Camel (Toma et al., 2023), along with studies on Flan-PaLM and PaLM-2’s medical reasoning using chain-of-thought (Singhal et al., 2023a, b; Wei et al., 2022; Wang et al., 2022).
+Researchers have also introduced Meditron (Chen et al., 2023b), an open-source medical LLM suite with 7B and 70B parameters, demonstrating superior performance against leading benchmarks and comparable results to GPT-3.5 and Med-PaLM-2.
+Despite the model size increase, performance gains have plateaued, and such expansions have not effectively addressed model forgetting (Wang et al., 2023; Luo et al., 2023).
+In response, our study integrates a retriever with LLMs during training to tackle this issue, offering 7B and 13B versions.
+Our findings show that these models exceed traditional pre-trained language models in performance metrics, with our 13B model surpassing even the Meditron 70B.
+Retrieval-Augmented Language Models
+In the clinical domain, LLMs’ factual inaccuracies due to gaps in medical knowledge pose significant risks, including misdiagnoses Petroni et al.
+(2019); Sung et al.
+(2021); Yao et al.
+(2022); Singhal et al.
+(2023a).
+Incorporating diverse knowledge repositories into LLMs enhances their performance across various NLP tasks, particularly in language modeling (Min et al., 2022).
+In traditional RAG models, the retrieval component is static, with updates applied only to the LLM.
+The process involves using the input as a query to first retrieve a collection of documents from a designated corpus, which the LLM then uses to enhance its predictions (Yu, 2022; Izacard et al., 2022).
+However, integrating retrieved documents does not guarantee effective use of external knowledge.
+Innovations like Atlas (Izacard et al., 2022) and RETRO (Borgeaud et al., 2022b) have modified models to better utilize retrieved information, but challenges remain in coherently blending this external data with the LLM’s pre-trained responses.
+Although some researchers have explored unsupervised training of retrievers to improve document relevancy for LLMs, such as REALM (Guu et al., 2020), REPLUG (Shi et al., 2023), and RA-DIT (Lin et al., 2023), these systems are usually trained independently from LLMs.
+Our approach innovatively integrates and synchronously trains both the retriever and the LLM, ensuring the retriever supplies more appropriate documents, thereby boosting the LLM’s effectiveness.
+We have validated our method’s superiority through comparisons with step-by-step training techniques, confirming its enhanced performance in seamlessly integrating external knowledge.
 9 Conclusion
@@
 Despite significant advancements in the performance of medical knowledge acquisition and question-answering systems, privacy protection becomes a paramount concern when dealing with sensitive medical data.
-In particular, our model utilizes data from textbooks and Uptodate.
-Textbooks are public datasets, while the UpToDate dataset cannot be publicly disclosed or used to produce derivative products based on their terms of use.
-We have sent an email to the official authorities and received their permission to use it.
-We would like to express our gratitude to UpToDate and must ensure that no information about UpToDate is disclosed.
+In particular, our model utilizes data from textbooks and public guideline.
 Bias Considerations
@@
 However, our model gave a more detailed and logical explanation of why there are decreased circulating T cells, as JLMR could retrieve related documents, such as some documents about SCID, to better understand the underlying principles.
-The last example differs from the first three, as the highest-scoring document retrieved was from UpToDate’s guideline.
+The last example differs from the first three, as the highest-scoring document retrieved was from public guideline.
 The guideline provides a detailed introduction to PCOS and its symptoms, helping the model better explain.
@@
 Therefore, careful monitoring and slow correction of hyponatremia are essential in this patient to avoid this potentially fatal complication.
-|
+| |
 | Our: This patient’s presentation is consistent with alcohol withdrawal syndrome (AWS), which is characterized by autonomic hyperactivity, tremulousness, and seizures.
@@
 Rapid correction of his sodium levels is not indicated due to the risk of osmotic myelinolysis.
-|
+| |
 | Expertise Comment: ChatGPT incorrectly interpreted the patient’s symptoms as primarily related to Alcohol Withdrawal Syndrome (AWS) without adequately considering the patient’s laboratory results, particularly the severe hyponatremia (serum sodium 112 mEq/L).
@@
 Termination of pregnancy (Option B) is not necessary unless there are severe fetal abnormalities detected on further evaluation.
-|
+| |
 | Our: The patient’s history of a rash that started on her face and progressed downwards, along with the presence of postauricular lymphadenopathy, is suggestive of rubella infection.
@@
 Intravenous immunoglobulin therapy is not indicated for rubella infection during pregnancy.
-|
+| |
 | Expertise Comment: ChatGPT made a made a basic error; it’s not appropriate to vaccinate during pregnancy. |
-| 1 | Q3: A 1-year-old girl is brought to the physician… |
+| 1 | Q4: A 1-year-old girl is brought to the physician… |
 | Answer: | |
@@
 Other options such as increased circulating IgE (Option B) or increased circulating neutrophils (Option F) are less specific to SCID and could be seen in a variety of other conditions.
-|
+| |
 | Our: Correct Answer: D Decreased circulating T cells This patient’s severe combined immunodeficiency (SCID) is most likely due to an autosomal recessive adenosine deaminase deficiency (the second most common cause of SCID), which results in disrupted purine metabolism and accumulation of metabolites (including deoxyadenosine and dATP) that are toxic to maturing lymphocytes.
@@
 Clinically, patients with SCID have no tonsils, no palpable lymph nodes, and a rudimentary thymus.
-In contrast, X-linked SCID (the most common variant) is caused by a defect in the IL-2 receptor, which is required for T |
+In contrast, X-linked SCID (the most common variant) is caused by a defect in the IL-2 receptor, which is required for T | |
 | Expertise Comment: ChatGPT’s explanation of the pathology is unclear, such as the specific type of SCID (Severe Combined Immunodeficiency caused by adenosine deaminase deficiency) and how this condition affects the proliferation and maturation of lymphocytes, leading to a reduced number of T cells.
@@
 Delayed umbilical cord separation would also not be expected… |
-| 1 | Q4: A 27-year-old woman presents to the office with concerns about her long struggle with her physical appearance since adolescence.
+| 1 | Q3: A 27-year-old woman presents to the office with concerns about her long struggle with her physical appearance since adolescence.
 She says she has always been l̈argeänd was constantly targeted by her classmates and coworkers for being so.
@@
 However, PCOS is a more common cause of hirsutism and irregular menstrual cycles in young women.
-|
+| |
 | Our: The patient’s history of being "large" since adolescence, a family history of type 2 diabetes mellitus, and a menstrual cycle that occurs every 45 days, along with her physical examination findings of severe inflammatory acne over the cheeks and forehead and dark coarse hairs on the back, are suggestive of polycystic ovary syndrome (PCOS).
@@
 E: Ovarian hyperthecosis is a condition characterized by excessive production of androgens by the ovaries, which can lead to hirsutism and acne, but it does not typically present with the other symptoms or findings seen in this patient.
-|
+| |
 | Expertise Comment: Our explanation is slightly more detailed in terms of linking the patient’s specific symptoms to the characteristics of PCOS and explaining why other diagnoses are less likely.
@@
 On the other hand, ChatGPT offers a more concise but still accurate diagnosis.
-| |
-Retrieve: The clinical features of PCOS are described here briefly but are reviewed in detail separately.
-(See "Clinical manifestations of polycystic ovary syndrome in adults".) PCOS is thought to be one of the most common endocrinopathies in women, affecting between 5 and 10 percent of women, depending upon the population studied [1].
-(see "Epidemiology, phenotype, and genetics of the polycystic ovary syndrome in adults", section on ’Epidemiology’).
-The syndrome is characterized clinically by oligomenorrhea and hyperandrogenism, as well as the frequent presence of associated risk factors for cardiovascular disease, including obesity, glucose intolerance, dyslipidemia, fatty liver, and obstructive sleep apnea.
-Other features include:… (Robert L Barbieri and William F Crowley, 2024) |
+|
+| MMLU | MedMCQA | MedQA | Amboss | Average | ||
+| Non Corpus | PMC-Llama-7B | 59.7 | 57.6 | 42.4 | 43.7 | 50.9 |
+| Llama 2-7B | 56.3 | 54.4 | 44.0 | 46.5 | 50.3 | |
+| Meditron-7B | 55.6 | 59.2 | 47.9 | 50.1 | 53.2 | |
+| ChatGPT | 69.4 | 51.0 | 50.2 | 49.1 | 54.9 | |
+| Meditron70B | 73.6 | 65.1 | 60.7 | 76.4 | 68.9 | |
+| Open Guidelines | RAG-7B | 63.8 | 62.1 | 47.3 | 50.7 | 55.9 |
+| RAG-13B | 69.8 | 63.4 | 56.8 | 60.7 | 62.7 | |
+| RAG-ChatGPT | 68.9 | 55.2 | 53.3 | 49.3 | 56.7 | |
+| JLMR-7B | 65.3 | 64.1 | 51.3 | 68.3 | 62.3 | |
+| JLMR-13B | 70.1 | 64.5 | 59.5 | 79.6 | 68.4 | |
+| All Corpus | RAG-7B | 62.1 | 62.4 | 54.6 | 70.7 | 62.5 |
+| RAG-13B | 69.9 | 64.2 | 59.9 | 76.9 | 67.7 | |
+| JLMR-ChatGPT | 70.1 | 55.3 | 54.3 | 50.1 | 57.5 | |
+| JLMR-7B | 64.3 | 62.4 | 56.2 | 71.2 | 63.5 | |
+| JLMR-13B | 72.8 | 65.5 | 62.5 | 81.2 | 70.5 | |
+| MedRAG Corpus | MedRAG-ChatGPT | 75.5 | 58.0 | 53.6 | 48.8 | 59.0 |